OpenStudy (anonymous):

Write the equation of a line in slope intercept form that is perpendicular to the line y = –4x and passes through the point (2, 6). Do not use any spaces. Slope should be a reduced improper fraction.

OpenStudy (anonymous):

first we have to find the perpendicular slope and we can do that by flipping the bottom number with the top number in the fraction 4/1 then multiplying it by -1  now that we have the slope we can plug the given imformation into this formula! \[y - y_{1} = m(x-x_{1})\] y1=6 x1=2 m=-1/4 now just solve for y and you will have your equation! \[y - 6 = -\frac{1}{4}(x-2)\]